OpenStudy (anonymous):

The area of a rectangle is 54 ft2 , and the length of the rectangle is 3 ft more than twice the width.

OpenStudy (anonymous):

@iambatman @iambatman @aum @dareintheren @PurplePanda312 @Race100

OpenStudy (anonymous):

Find the dimensions of the rectangle.

OpenStudy (anonymous):

@theEric @tkhunny @Teddyiswatshecallsme @Rockiee @girlover @ash2326

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@PurplePanda312

OpenStudy (anonymous):

I don't know sorry.

OpenStudy (anonymous):

@hartnn @ash2326

OpenStudy (anonymous):

@fateal

OpenStudy (fateal):

lol call @Dark_Riddles this is a goo riddle :')

OpenStudy (anonymous):

@Dark_Riddles

OpenStudy (anonymous):

@Dark_Riddles

OpenStudy (anonymous):

c

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

The area of a rectangle is 54 ft2 , and the length of the rectangle is 3 ft more than twice the width.

OpenStudy (anonymous):

Find the dimensions of the rectangle.

OpenStudy (anonymous):

Well, I'll tell you one thing, this house is going to be crooked as crap.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how do I put this into an equation

OpenStudy (anonymous):

are you confuse as I

OpenStudy (anonymous):

is this right X^2+3x-54=0 @Dark_Riddles

OpenStudy (tkhunny):

Well, using "X" is probably a bad idea. Changing to "x" int eh middle of the problem is another bad idea. Super Big Clue that will Change Your Life --- WRITE DOWN DEFINITIONS!! L = Length W = Width L*W = Area "area of a rectangle is 54 ft2" L*W = 54 " length of the rectangle is 3 ft more than twice the width." L = 3 + 2W Substitute for L (3+2W)*W = 54 2W^2 + 3W - 54 = 0 Notice how every step of the way you can FOLLOW what is happening. This is what you should be doing.